OpenStudy (anonymous):

The ratio of the measures of two supplementary angles is 7:3. Find the measures of the angles. A. 126° and 54° B. 72° and 18° C. 159° and 21° D. 69° and 21°

jimthompson5910 (jim_thompson5910):

Hint: If "The ratio of the measures of two supplementary angles is 7:3", then 7x+3x = 180 Solve for x. Then use this to find 7x and 3x

OpenStudy (anonymous):

10x=180

OpenStudy (anonymous):

simplify

OpenStudy (anonymous):

i'm so confused !

jimthompson5910 (jim_thompson5910):

What is 180 divided by 10

OpenStudy (anonymous):

18

jimthompson5910 (jim_thompson5910):

So x = 18

jimthompson5910 (jim_thompson5910):

Now use this to find 7x and 3x

OpenStudy (anonymous):

Right so plug in 18 as x

jimthompson5910 (jim_thompson5910):

I'll do 7x 7x = 7*18 7x = 126

OpenStudy (anonymous):

7*18 and 3*18=180

jimthompson5910 (jim_thompson5910):

3*18 = 54, not 180

OpenStudy (anonymous):

You just gave her the answer lol. i put everything in one equation so she could see it better...

jimthompson5910 (jim_thompson5910):

oh i see what you did, you just wrote it in an odd way

OpenStudy (anonymous):

lol sorry

jimthompson5910 (jim_thompson5910):

that's ok

OpenStudy (anonymous):

lol its all good both of you , i appreciate the help ! thank you very much !!

OpenStudy (anonymous):

:)